Assume you just bought a new swimming pool that is circular and has a diameter of 15 feet and a height of 5 feet. How many gallons will it hold?
1
Expert's answer
2011-07-01T10:50:15-0400
The volume is the swimming pool is
V = Sh = π d2 h /4 = π*152*5/4 = 883.58 ft3.
In gallons 883.58/0.133 = 6643.41 gal.